If ages of sita gita are in the ratio 2/3 . after 3 years ,their ages are in the ratio 7/9. find their present ages.

Grade:7

1 Answers

Arun
5 years ago
Let age of sita and gita are of s and g years
 
s/g = 2/3
Means 3s = 2g
And
s+3/(g+3) = 7/9
 
9s + 27 = 7g + 21
 
g = 6
And s = 4 years
 
Regards
Arun (askIITians forum expert)